# Does Training on Synthetic Data Make Models Less Robust?

### v1

@@ Abstract
An increasingly common practice is to train large language models (LLMs) using synthetic data. Often this synthetic data is produced by the same or similar LLMs as those it is being used to train. This raises the question of whether the synthetic data might in fact exacerbate certain “blindspots” by reinforcing heuristics that the LLM already encodes. In this paper, we conduct simulated experiments on the natural language inference (NLI) task with Llama-2-7B-hf models. We use MultiNLI as the general task and HANS, a targeted evaluation set designed to measure the presence of specific heuristic strategies for NLI, as our “blindspot” task. Our goal is to determine whether performance disparities between the general and blind spot tasks emerge. Our results indicate that synthetic data does not reinforce blindspots in the way we expected. Specifically, we see that, while fine-tuning with synthetic data doesn’t necessarily reduce the use of the heuristic, it also does not make it worse as we hypothesized.
@@ 1 Introduction and Related Work
In this work, we hypothesize that, because synthetic data less diverse than the original training data Whitney and Norman (2024), it is more likely to have blindspots and thus that fine-tuning on model-generated data will exacerbate these blindspots in the tuned model. In particular, we hypothesize that the synthetic data will encode the heuristic to a larger extent that would naturally occurring data, and thus that fine-tuning on synthetic data will lead the model to more strongly favor the heuristic. This weakness would be revealed in data that specifically is designed to test whether models are using the heuristic, as models trained on synthetic data might still show improved performance on generic test sets on which the heuristic performs well.
@@ 4 Conclusion
Another notable point is that pretrained large language models, such as Llama, encode a wealth of world knowledge. Many potential biases may have been addressed in the training process. On the other hand, human created or audited data are not inherently free from implicit biases. The more concentrated distribution and reduced diversity of synthetic data might reinforce biases in certain blindspot scenarios. However, the rich world knowledge embedded in the generator model can also address some biases, potentially outperforming humans in certain cases and contributing positively to bias mitigation. A critical direction for future research is to disentangle these two effects and assess the significance of each, thereby enhancing our understanding of the impact of training models with synthetic data.

### v2

@@ Abstract
An increasingly common practice is to train large language models (LLMs) using synthetic data. Often this synthetic data is produced by the same or similar LLMs as those it is being used to train. This raises the question of whether the synthetic data might in fact exacerbate certain “blindspots” by reinforcing heuristics that the LLM already encodes. In this paper, we conduct simulated experiments on the natural language inference (NLI) task with Llama-2-7B-hf models. We use MultiNLI as the general task and HANS, a targeted evaluation set designed to measure the presence of specific heuristic strategies for NLI, as our “blindspot” task. Our goal is to determine whether performance disparities between the general and blind spot tasks emerge. Our results indicate that synthetic data does not reinforce blindspots in the way we expected. Specifically, we see that, while fine-tuning with synthetic data doesn’t necessarily reduce the use of the heuristic, it also does not make it worse as we hypothesized.  [^1] Our code is available at https://github.com/untakenJ/synthetic-data-blindspot.
@@ 1 Introduction and Related Work
In this work, we hypothesize that, because synthetic data less diverse than the original training data Whitney and Norman (2024), it is more likely to have blindspots and thus that fine-tuning on model-generated data will exacerbate these blindspots in the tuned model. In particular, we hypothesize that the synthetic data will encode the heuristic to a larger extent than naturally occurring data would, and thus that fine-tuning on synthetic data will lead the model to more strongly favor the heuristic. This weakness would be revealed in data that specifically is designed to test whether models are using the heuristic, as models trained on synthetic data might still show improved performance on generic test sets on which the heuristic performs well.
@@ 4 Conclusion
Another notable point is that pretrained large language models, such as Llama, encode a wealth of world knowledge. Many potential biases may have been mitigated during training, particularly for models deployed in real-world applications, which are typically much larger and more powerful than the fine-tuned Llama-2-7B-hf generator used in our study. On the other hand, human-created or audited data are not inherently free from implicit biases. The more concentrated distribution and reduced diversity of synthetic data might reinforce biases in certain blindspot scenarios. However, the rich world knowledge embedded in the generator model can also address some biases, potentially outperforming humans in certain cases and contributing positively to bias mitigation. A critical direction for future research is to disentangle these two effects and assess the significance of each, thereby enhancing our understanding of the impact of training models with synthetic data.